### 23 Mar 2026 v3

# Thibaud Cambronne12, Samuel Bobick13, Wente Zeng4, Scott Moura12

Abstract
Demand charge, a utility fee based on an electricity customer’s peak power consumption, often constitutes a significant portion of costs for commercial electric vehicle (EV) charging station operators. This paper explores control methods to reduce peak power consumption at workplace EV charging stations in a joint price and power optimization framework. We optimize a menu of price options to incentivize users to select controllable charging service. Using this framework, we propose a model predictive control approach to reduce both demand charge and overall operator costs. Through a Monte Carlo simulation, we find that our algorithm outperforms a state-of-the-art benchmark optimization strategy and can significantly reduce station operator costs.
I Introduction
Smart electric vehicle (EV) charging at workplaces presents many key opportunities to contribute to transportation electrification and decarbonization. For example, although EV owners typically charge overnight, midday workplace charging capitalizes on cheap, renewable-laden electricity, thus realizing the full emissions reduction potential of EVs [powell2022charging]. Also, workplace charging can alleviate inequities for EV owners without access to at-home chargers, such as those in multi-unit dwellings [peterson2011addressing]. Furthermore, intelligent charging algorithms can help align charging with utility time-of-use (TOU) rate schedules and shave peak power consumption to maximize profit and incentivize charger expansion by private companies [williams2014pricing].
A major cost for some EV charging station operators is demand charge, a utility fee based on the highest power consumption during a billing cycle. Approximately 5 million commercial customers in the United States are estimated to face retail electricity tariffs with demand charges greater than \$15/kW, accounting for more than a quarter of the nation’s 18 million commercial customers [mclaren2017identifying]. Demand charges are a significant component of commercial utility bills, generally accounting for 30% to 70% of total electricity costs [mclaren2017identifying].
Demand charge reduction through smart charging presents a challenging control problem. Even with smart charging technology, EV charging station arrivals are stochastic. On busy days, a single ill-timed arrival can push the station’s load above the running peak, increasing demand charge.
Most prior work on demand charge mitigation for EV charging stations focuses on optimizing power schedules to maximize profit while accounting for EV arrival and departure times, energy demand constraints, demand charges, and TOU utility tariffs [lee2020pricing, powell-transformer, kara-benefits, chaudhari2022learning, Yang2024]. Lee, Pang, and Low [lee2020pricing] propose an offline pricing scheme that optimizes power delivery under fixed arrival, departure, and energy demand constraints. Their approach ensures cost recovery by incorporating facility demand charges, time-varying energy costs, and congestion costs. However, it does not account for stochastic variations in EV charging behavior. Yang et al. [Yang2024] use “block” model predictive control to optimize power delivery for demand charge mitigation under uncertain future arrivals, departures, and energy demands. However, their approach does not consider how to structure pricing to encourage user behavior that mitigates demand charge.
Formulations in [zeng2021, obeid2023learning, Ozturk2025] optimize both power delivery and user-facing prices. By incentivizing users with discounts if they allow their load to be controlled instead of uncontrolled, they are able to increase revenue. However, these formulations only control on-going charging sessions. At workplace stations, most charging sessions begin in the morning and continue until the workday ends. Consequently, control decisions made at 8 AM can influence peak power at 2 PM. Therefore, early indicators of a peak-inducing day should be incorporated into morning control decisions. An effective controller needs anticipatory ability.
To solve this problem, we propose two solutions that provide anticipatory abilities: (i) convex reformulations of the demand charge constraint, and (ii) a forecast-enabled model predictive control (MPC) approach. This enables the following novel contributions to the existing literature:
  1. 1.
    We propose a MPC method that anticipates demand spikes using time series forecasting.
  2. 2.
    We increase the load flexibility available to the controller by leveraging dynamic pricing to incentivize users to choose a controllable charging service.
  3. 3.
    We demonstrate that our approach outperforms the benchmark optimization algorithm from [zeng2021] and other non-MPC reformulations in reducing both demand charge and overall operator costs.
II Benchmark Optimization Problem
In this section, we present a benchmark station-level optimization problem that is solved when a new user arrives at the charging station. Our formulation closely follows the formulation presented in [zeng2021], with some modifications.
Upon arrival at the charging station, user $i$ inputs their required energy ${\text{E}}_{i}^{\text{req}}$ and expected departure time from the charging station $T_{i}^{\text{end}}$. Given this information, the user is presented with two charging prices, $z_{\text{sch}}$ and $z_{\text{reg}}$, generated by a price and power optimization algorithm. Given these prices, the user can make one of three choices:
  1. 1.
    REGULAR: the battery charges at maximum power until it tops off or unplugs. Importantly, this load is uncontrollable.
  2. 2.
    SCHEDULED: the user’s energy required is guaranteed to be provided by their indicated time of departure. The power delivery is scheduled with a control algorithm.
  3. 3.
    LEAVE: The user can leave if they consider the prices to be too high, opting for nearby chargers or a parking spot without a charger. Without this option, the optimal service price that maximizes profit is infinite, which is non-sensical.
The control objective is to maximize the expected profit from the station by choosing the REGULAR and SCHEDULED prices presented to the new user, along with the power profiles for all users who selected the SCHEDULED charging option.
The benchmark solution is a conventional non-anticipatory optimization that involves solving (1) directly. We assume that users choose one of the 3 charging options based solely on the price menu. We denote $\text{Pr}\left(m\mid z\right)$ to be the probability that new user $n$ chooses option $m$ when presented with the set of prices $z$. We model these probabilities with a two-step discrete choice model, described in Appendix A. Due to the non-convexity of the discrete choice model (A.5)–(A.9), optimization problem (1) is difficult to solve numerically. As such, we find an approximate solution to (1) by searching over a grid of candidate prices $z_{\text{candidate}}=(z_{\text{sch}},z_{\text{reg}})$. At each point in the grid, we solve $\min_{p}\mathsf{E}[{f(p,m)\mid z_{\text{candidate}}}]$, which is a linear program.
II-A Objective Function
The joint price and power objective, defined as a function of prices $z$ and power profiles $p$, is expressed as an expectation over new user’s choice $m$. We formulate the resulting optimization problem as follows:
$\min_{z,p}$~~ $\mathsf{E}[f(p,m)\mid z]$ (1)
$=$ $\ \text{Pr}\left(m=\textit{SCHEDULED}{}\mid z\right)f^{\text{sch}}(z,p)$
$+$ $\ \text{Pr}\left(m=\textit{REGULAR}{}\mid z\right)f^{\text{reg}}(z,p)$
$+$ $\ \text{Pr}\left(m=\textit{LEAVE}{}\mid z\right)f^{\text{leave}}$
[…]
Here, $f^{\text{sch}}$, $f^{\text{reg}}$, and $f^{\text{leave}}$ represent the station operator’s cost functions associated with each potential choice made by the new user. The optimization horizon runs from current time index $T^{\text{start}}$ up to but not including $T^{\text{end}}$, the latest indicated time of departure across all users $i\in\mathcal{I}$.
II-B Demand Charge Constraints
The running peak power is tracked over a monthly billing cycle, consistent with utility practice. The billing cycle starts at $\tau=0$. $G_{\tau}$ is the station load at time $\tau$:
$G_{\tau}$ $=\sum_{i\in A_{\text{sch}}}p_{i,\tau}^{\text{sch}}+\sum_{j\in A_{\text{reg}}}p^{\text{max}}\mathsf{1}\{\tau<T_{j}^{\text{end}}\}\quad\forall\tau\geq 0.$ (2)
The running peak power $D_{\tau}$ is defined as follows:
$D_{0}$ $=0,$ (3)
$D_{\tau+1}$ $=\max\{G_{\tau+1},D_{\tau}\}\quad\forall\tau\geq 0.$ (4)
II-C Scheduled Cost Function
Provided with new user $n$’s energy required $E_{n}^{\text{req}}$ and time index of departure $T_{n}^{\text{end}}$, the cost function assuming that user $n$ chooses the SCHEDULED charging option is given by
$f^{\text{sch}}\left(z,p\right)$ $=\sum_{\tau=T^{\text{start}}}^{T_{n}^{\text{end}}-1}\left(c_{\tau}-z^{\text{sch}}\right)p_{n,\tau}^{\text{sch}}\Delta t$ (5)
$+\sum_{i\in\mathcal{A}_{\text{sch}}}\left[\sum_{\tau=T^{\text{start}}}^{T^{\text{end}}_{i}-1}\left(c_{\tau}-\zeta_{i}\right)p_{i,\tau}^{\text{sch}}\Delta t\right]$
$+\sum_{j\in\mathcal{A}_{\text{reg}}}\left[\sum_{\tau=T^{\text{start}}}^{T^{\text{end}}_{j}-1}\left(c_{\tau}-\zeta_{j}\right)p^{\text{max}}\Delta t\right]$
$+c_{D}\left(D_{T^{\text{end}}-1}-D_{T^{\text{start}}-1}\right).$
The final term of the objective, $c_{D}\left(D_{T^{\text{end}}-1}-D_{T^{\text{start}}-1}\right)$, represents the cost associated with the increase in monthly peak power over the optimization horizon $T^{\text{start}},T^{\text{end}})$. Note that, as indicated by (4), this term cannot be negative and only becomes positive when the monthly peak power increases over the optimization horizon.
II-D Regular Cost Function
Provided with new user $n$’s energy required $E_{n}^{\text{req}}$ and the time index $T_{n}^{\text{end}}$ when the user’s battery would be fully charged when charging at $p^{\text{max}}$, the cost function assuming that user $n$ chooses the REGULAR charging option is given by
$f^{\text{reg}}\left(z,p\right)$ $=\sum_{\tau=T^{\text{start}}}^{T_{n}^{\text{end}}-1}\left(c_{\tau}-z^{\text{reg}}\right)p^{\text{max}}\Delta t$ (6)
$+\sum_{i\in\mathcal{A}_{\text{sch}}}\left[\sum_{\tau=T^{\text{start}}}^{T^{\text{end}}_{i}-1}\left(c_{\tau}-\zeta_{i}\right)p_{i,\tau}^{\text{sch}}\Delta t\right]$
$+\sum_{j\in\mathcal{A}_{\text{reg}}}\left[\sum_{\tau=T^{\text{start}}}^{T^{\text{end}}_{j}-1}\left(c_{\tau}-\zeta_{j}\right)p^{\text{max}}\Delta t\right]$
$+c_{D}\left(D_{T_{\text{end}}-1}-D_{T_{\text{start}}-1}\right).$
II-E Leave Cost Function
If the user chooses to leave, the station operator incurs neither cost nor benefit. Thus, $f^{\text{leave}}=0$.
II-F Energy Constraints
Constraint (7) ensures that the charging power lies within the charger rating, (8) tracks the battery state of charge $e_{i,\tau}$, and (9) ensures that the chosen power profiles satisfy each user $i\in\mathcal{I}$’s energy requirements $E_{i}^{\text{req}}$:
$0$ $\leq p_{i,\tau}^{\text{sch}}\leq p^{\text{max}}$ $\forall i\in\mathcal{I},\;\tau\in[T^{\text{start}},T_{i}^{\text{end}}),$ (7)
$e_{i,\tau}$ $=e_{i,\tau-1}+\Delta t\eta p_{i,\tau}$ $\forall i\in\mathcal{I},\;\tau\in[T^{\text{start}},T_{i}^{\text{end}}),$ (8)
$E_{i}^{\text{req}}$ $\leq e_{i,T_{i}^{\text{end}}}$ $\forall i\in\mathcal{I}.$ (9)
III Controllers for Improved Peak Power Management
In this section, we propose several modifications and improvements to optimization problem (1) which aim to reduce monthly demand charge. Section III-A applies a hard constraint to peak power. Section III-B gives extra weight to the demand charge term when the current power profile is on the verge of exceeding peak power. Section III-C uses MPC to give the controller an anticipatory ability.
III-A Iterative Hard Thresholding
In this approach, we add a hard constraint $M$ on station load $G_{\tau}$ to find the feasible solution that results in the lowest increase in peak power. That is, we solve (1) with the additional constraint
$G_{\tau}\leq M\quad\forall\tau\in[T^{\text{start}},T^{\text{end}}).$ (10)
$M$ is initialized at the running peak in the billing cycle so far, $D_{T^{\text{start}}-1}$. If (1) with constraint (10) is infeasible, we iteratively loosen the constraint by increasing $M$ by some step size $\epsilon>0$, and re-solve until we find a feasible solution.
III-B Softplus Demand Charge Penalty
Approaching the running peak is undesirable because any future arrival is likely to increase peak power. Thus, we penalize this case by applying the convex softplus function, $\text{softplus}(x)=\log(1+e^{x})$, to the increase in peak power over the billing cycle, as illustrated in Fig. 1. That is, we replace the demand charge terms in (5) and (6) with
$c_{D}\text{softplus}\left(D_{T^{\text{end}}-1}-D_{T^{\text{start}}-1}\right).$ (11)
Figure 1: The softplus function adds a penalty when the peak of the optimized station power is approaching the running peak (as $x\xrightarrow{}0^{-}$), and converges to the original penalty as the station power exceeds the running peak (as $x\xrightarrow{}+\infty$).
III-C Model Predictive Control
Next, we incorporate a time series forecast into the demand charge term to anticipate arrivals. Let $\{\hat{G}_{T_{\text{start}}\mid m},...,\hat{G}_{T_{\text{start}}+l\mid m}\}$ represent the forecasted station power time series for the next $l$ timesteps, given that new user $n$ chooses charging option $m$. Then, we can replace the demand charge terms in (5) and (6) with
$c_{D}\left(\max\{\hat{G}_{T^{\text{start}}\mid m},...,\hat{G}_{T^{\text{start}}+l\mid m},D_{T^{\text{start}}-1}\}-D_{T^{\text{start}}-1}\right).$ (12)
We test the MPC algorithm with 3 types of forecasters:
  1. 1.
    A naive model that assumes no additional arrivals, but uses, as a forecast, the sum of the optimal power profile from the previous optimization with new user $n$’s power profile, assuming they chose the REGULAR option;
  2. 2.
    A linear model trained with elastic net regression [zou2005regularization];
  3. 3.
    XGBoost [chen2016xgboost].
Forecasting models 2) and 3) use features such as recent historical power profiles, planned power profiles from the last optimization, number of sessions, and encodings of time of day, week, year, and workday. Model hyperparameters are tuned with a train-validation-test split of empirical station usage data. More details on feature engineering and forecast implementation can be found in Appendix B.
IV Numerical Study
TABLE I: Profit Comparison Across Control Schemes
Control Scheme Mean Cost or Revenue (\$) Change from Benchmark (%)
Demand Charge TOU Revenue Cost Profit Demand Charge TOU Cost
Benchmark 624 1,984 4,309 2,608 1,702 0.00 0.00 0.00
Threshold 628 1,987 4,331 2,615 1,716 -0.71 0.16 +0.29
Softplus 609 2,017 4,446 2,626 1,819 -2.34 1.67 +0.71
MPC (Naive) 519 1,970 4,253 2,488 1,765 -16.90 -0.71 -4.58
MPC (Linear) 521 1,970 4,362 2,491 1,870 -16.50 -0.68 -4.47
MPC (XGBoost) 533 1,956 4,434 2,489 1,944 -14.57 -1.39 -4.55
IV-A Simulation Overview
We perform a Monte Carlo simulation [^1] Simulation code available at https://github.com/samuelBobick/StationLevelPowerForecasting. to quantitatively validate the performance of the control algorithms presented in the previous section. We replay all EV charging sessions sessions on the Smart Learning Pilot for EV Charging Stations (SlrpEV) platform from 2023, a total of 2274 charging sessions [Ozturk2025]. We assume a demand charge of \$20/kW [mclaren2017identifying], representing 24% of simulated total station operator costs under the benchmark algorithm. We use the Pacific Gas & Electric Commercial Business Electric Vehicle June 2023 rates for the TOU tariff structure [pge_tariff]. Note that the computational complexity of the algorithms proposed in Section III is not prohibitive; all algorithms run within in a fraction of a second on a standard laptop. Further details on the simulation setup can be found in Appendix C.
IV-B Results and Discussion
Table I presents the Monte Carlo simulation results, obtained by running each of the 12 months 10 times.
In particular, all MPC schemes successfully decrease demand charge by approximately 15% from the benchmark, while also decreasing TOU costs by approximately 1%, resulting in an approximately 4.5% average total cost reduction. The two non-MPC controllers fail to significantly reduce the total costs.
Figure 2 demonstrates how MPC improves peak power management. As the day’s first sessions begin, the benchmark controller waits to fulfill energy demand until the cheapest energy (super off-peak) is available. On the other hand, the MPC controller anticipates a busy day, and starts charging earlier so that there is less demand during the peak power event. In fact, by aggressively charging early in the day, the MPC algorithm is even able to partially avoid peak TOU pricing periods.
Figure 2: Example of control actions for benchmark (left) and MPC (right) on September 6th, 2023. For these two examples, the users all arrived at the same time, with the same requirements and all chose SCHEDULED. Each color represents a single user’s power profile. The MPC algorithm shows better repartition of the load throughout the off peak hours and achieves a lower peak (24.57 kW) compared to the benchmark solution (28.64 kW), while also reducing the energy consumed during peak TOU hours.
Figure 3 illustrates how the MPC controller leverages pricing to encourage users to select SCHEDULED when demand charge mitigation is pertinent. Both strategies raise prices before and during peak TOU hours when energy costs double. However, the MPC algorithm sets high prices early in the morning while significantly discounting SCHEDULED between 6–8 AM, prioritizing a base of controllable SCHEDULED sessions for the day. As a result, the MPC controller is able to maintain lower daytime prices, particularly for REGULAR sessions, whereas the benchmark controller must raise prices during peak load hours to compensate for unanticipated sessions.
Figure 3: Example of pricing strategy for benchmark (top) and MPC (bottom) solutions, overlaid with the average station power profile.
IV-C Forecast Performance vs. MPC Peak Power Reduction
Table II compares forecast errors to simulated mean peak power. The training RMSE closely correlates with the RMSE of predictions generated during the simulation. However, the naive forecast achieves the best peak power reduction despite a much higher training RMSE and a higher simulation RMSE. This suggests that minimizing RMSE alone does not strongly correspond to improved MPC performance.
TABLE II: The Value of Forecast Accuracy in MPC
Control Scheme Training RMSE (kW) Simulation RMSE (kW) Mean Peak Power (kW)
MPC (Naive) 11.36 7.46 25.93
MPC (Linear) 6.25 4.13 26.05
MPC (XGBoost) 5.94 4.31 26.65
One possible explanation for this is that the linear and XGBoost models were trained on data that does not exactly match the distribution of simulation data. Specifically, the training data contains charging sessions which were controlled with an optimization algorithm similar to the benchmark algorithm. On the other hand, the naive model is based on the previous output of the MPC controller, making it a perfect forecast if no new vehicles arrive on that day and a good approximation even if a few do. Therefore, retraining the models on data generated by the MPC controller would better match forecasts with their role in the optimization formulation. But this is challenging, as forecasts depend on MPC outputs cannot be determined without the forecasts, creating a self-referential problem.
Our test case, the SlrpEV station, has eight chargers. Due to its small size, station loads are stepwise and irregular, making forecasting more challenging. For a larger set of EV chargers, the Law of Large Numbers would smooth load curves, increasing periodicity and improving forecast accuracy. Consequently, the MPC approach may perform even better at larger stations or across a collection of aggregated station loads.
V Conclusions and Limitations
In this paper, we present several methods to decrease demand charge at workplace EV charging stations in a joint price and power optimization scheme and use a Monte Carlo simulation to validate their performance against a state-of-the-art benchmark optimization algorithm. The MPC implementation detailed in Section III-C decreases overall station operator costs by 4.6% and demand charge by 16.9%.
In Section IV-C, we find that minimizing forecast RMSE alone does not strongly correlate with improved MPC performance. Li, Ju, and Wang [li2024quantifying] have a similar finding in the context of building energy management. Training the models on data generated by the MPC algorithm would be more accurate, but this introduces a self-referential problem that requires future work.
When simulating SlrpEV user sessions, we randomly select between REGULAR and SCHEDULED choices with probabilities specified by the discrete choice model. While we account for LEAVE in the optimization problem presented in (1), in the Monte Carlo simulation we are limited in the fact that users have already decided not to LEAVE. This limitation arises because the SlrpEV dataset only contains information about users who chose to charge. Thus, profit and revenue should be interpreted with caution. Ultimately, station operators optimize for profit, so further work is needed to incorporate LEAVE behavior into the Monte Carlo simulation.
Joint price and power optimization is both technically feasible and shown to have high user acceptance. Indeed, cloud-connected Level 2 power management is already offered by ChargePoint and EVgo, albeit without joint price and power optimization [chargepoint, evgo]. Moreover, prior work uses real-world experiments to show that significant proportions of customers are willing adopt SCHEDULED charging in exchange for discounts [obeid2023learning, Ozturk2025]. One shortcoming of our dataset is the fact that the SlrpEV station only has eight chargers. Future work should examine the scalability of joint price and power optimization considering peak power mitigation over tens, hundreds, or thousands of EV chargers by a single operator across many sites.
Ultimately, we find that MPC significantly decreases charging station operating costs, particularly demand charge. These savings can either boost operator profits or be used to offer lower prices to customers. Enhancing behavioral simulation, testing our algorithms for control of large aggregate loads, and experimentally deploying our control schemes with real human behavior as in [Ozturk2025, lee-adaptive] are key steps toward further improvement of joint price and power control for demand charge mitigation at workplace EV charging stations.
Appendix A Behavioral Model Formulation
Our behavioral model describes the probability that user $n$ selects charging choice $m$ when presented with prices $z$. Since $z$ represents prices per unit of energy (\$/kWh), we multiply it by the maximum charging power $p^{\text{max}}$ to convert it to a price per unit of time (\$/hr). This transformation ensures consistency between the optimization framework in Section II and the utility model in [Ozturk2025].
$z_{\text{reg}}^{\prime}=z^{\text{reg}}p^{\text{max}},\quad z_{\text{sch}}^{\prime}=z^{\text{sch}}p^{\text{max}}$ (A.1)
To estimate the choice probabilities, we use a two-step discrete choice model, using utilities $U$ estimated in [Ozturk2025].
$U_{\text{reg}}$ $=0.341-0.0184\frac{(z^{\prime}_{\text{reg}}-z^{\prime}_{\text{sch}})}{2}$ (A.2)
$U_{\text{sch}}$ $=0.0184\frac{(z^{\prime}_{\text{reg}}-z^{\prime}_{\text{sch}})}{2}$ (A.3)
$U_{\text{leave}}$ $=-1+0.005\frac{(z^{\prime}_{\text{reg}}+z^{\prime}_{\text{sch}})}{2}$ (A.4)
First, we calculate the probability that the user chooses the LEAVE option:
$\text{Pr}\left(m=\textit{LEAVE}{}\mid z\right)$ $=\frac{e^{U_{\text{leave}}}}{e^{U_{\text{sch}}}+e^{U_{\text{reg}}}+e^{U_{\text{leave}}}}.$ (A.5)
Then, we calculate the probability that the user chooses SCHEDULED, given that they do not choose the LEAVE option:
$\text{Pr}(m$ $=\textit{SCHEDULED}{}\mid z,m\neq\textit{LEAVE}{})$ (A.6)
$=\frac{e^{U_{\text{sch}}}}{e^{U_{\text{sch}}}+e^{U_{\text{reg}}}}(1-\text{Pr}(m=\textit{LEAVE}{}\mid z)).$ (A.7)
Likewise, for REGULAR,
$\text{Pr}(m$ $=\textit{REGULAR}{}\mid z,m\neq\textit{LEAVE}{})$ (A.8)
$=\frac{e^{U_{\text{reg}}}}{e^{U_{\text{sch}}}+e^{U_{\text{reg}}}}(1-\text{Pr}(m=\textit{LEAVE}{}\mid z)).$ (A.9)
Appendix B Time Series Forecasting Details
Our model forecasts the station power time series for the next $l$ timesteps. It takes the following as features [^2] More external features (such as weather, special events, etc.) could be further introduced to improve the prediction accuracy. However, we don’t consider them in this paper, as the focus of this work is the MPC formulation.:
  1. 1.
    The $k$ most recent aggregate power values: $G_{T_{\text{start}}-k},G_{T_{\text{start}}-k+1},...G_{T_{\text{start}}-1}$;
  2. 2.
    The planned station power profile for the next $l$ timesteps for all active users in $\mathcal{I}$, considering that user $n$ chooses REGULAR, and given power profiles $\bar{p}$ generated from the previous optimization (triggered when user n1n-1 arrived) for users in $\mathcal{A}_{\text{sch}}\cup\mathcal{A}_{\text{reg}}$: $G_{T_{\text{start}}|\textit{REGULAR}{},\bar{p}},...G_{T_{\text{start}}+l|\textit{REGULAR}{},\bar{p}}$;
  3. 3.
    The number of active sessions at the charging station, including new user $n$: $|\mathcal{I}|$;
  4. 4.
    The total number of working chargers at the station at time index $T^{\text{start}}$;
  5. 5.
    A sinusoidal positional encoding of the time of day, time of week, and time of year [vaswani2017attention];
  6. 6.
    A boolean variable that evaluates to 1 if the time index $T^{\text{start}}+24$ hours falls on a workday.
Let $\psi_{\tau}$ represent the forecast of the station power profile, made outside of the control loop. At each solver iteration, to calculate $\hat{G}_{\tau\mid\textit{REGULAR}{}}$ and $\hat{G}_{\tau\mid\textit{SCHEDULED}{}}$ for time indices $\tau\in[T^{\text{start}},T^{\text{end}})$, we can simply start from the original forecast $\psi_{\tau}$ and add and subtract power profiles of the new control input $p$. Let $\bar{p}^{\text{sch}}_{i}$ represent the optimal scheduled power profile for user $i$ found in the previous optimization triggered by user n1n-1. Then, we update our forecast as follows:
$\hat{G}_{\tau\mid\textit{REGULAR}{}}=\psi_{\tau}+\sum_{i\in\mathcal{A}_{\text{sch}}}\left(p_{i,\tau}^{\text{sch}}-\bar{p}_{i,\tau}^{\text{sch}}\right)$ (A.10)
$\hat{G}_{\tau\mid\textit{SCHEDULED}{}}=\psi_{\tau}+(p_{n,\tau}^{\text{sch}}-p_{n,\tau}^{\text{reg}})+\sum_{i\in\mathcal{A}_{\text{sch}}}\left(p_{i,\tau}^{\text{sch}}-\bar{p}_{i,\tau}^{\text{sch}}\right)$ (A.11)
For each new user, the forecast is only executed once outside of the control loop, and is arithmetically updated to the new control actions at each grid search iteration of the solver. Compared to executing the forecast inside the control loop, our single forecast approach allows us to:
  1. 1.
    Improve forecaster performance by using non-convex forecasters versus being limited to convex models, or having to increase computational time by using non-convex optimization;
  2. 2.
    Improve the forecast reliability by making predictions only on controlled historical power profiles versus candidate power profiles tested by the solver, which may not match the distribution of data that the forecaster was trained on;
  3. 3.
    Reduce computational time by making a single prediction instead of one per solver iteration.
We employ separate models for workdays and non-workdays (i.e., weekends and holidays). The model is trained on the SlrpEV dataset [Ozturk2025]. Note that we post-process the forecast by clipping it to eliminate unrealistic forecasts:
$\psi_{\tau}$ $\geq p_{n,\tau}^{\text{reg}}+\sum_{i\in\mathcal{A}_{\text{sch}}}\bar{p}_{i,\tau}^{\text{sch}}+\sum_{i\in\mathcal{A}_{\text{reg}}}\bar{p}_{i,\tau}^{\text{reg}}$ (A.12)
$\forall\tau\in[T^{\text{start}},T^{\text{end}}).$
Appendix C Simulation Setup
The SlrpEV dataset used to derive our simulation has a mix of REGULAR and SCHEDULED charging sessions. As we re-generate user choices using the discrete choice model, we make several assumptions:
  1. 1.
    Users have already made the decision to charge. That is, they cannot LEAVE. They only choose between REGULAR and SCHEDULED options.
  2. 2.
    The energy demand for charging sessions that were in reality REGULAR but are randomly simulated to be SCHEDULED have an energy demand equal to 57% of the energy delivered in the original charging session, corresponding to the average in the SlrpEV dataset.
  3. 3.
    We assume we know that user $n$’s time of departure is fixed and known.
Additional parameter settings are listed in Table III.
TABLE III: Simulation Parameters
Parameter Value Description
$\Delta t$ 0.25 hours Length of each time step
$p^{\text{max}}$ 6.6 kW Maximum charging power
$\eta$ 1 Charger efficiency
$\epsilon$ 1 kW Threshold increment for solution III-A
$k$ 96 Number of past time steps to use as features for the forecaster
$l$ 32 Number of time steps ahead that the forecaster predicts

### 11 Apr 2026 v4

@@
 Through a Monte Carlo simulation, we find that our algorithm outperforms a state-of-the-art benchmark optimization strategy and can significantly reduce station operator costs.
+Nomenclature
+  • Indices/Sets
+  • $\tau/\mathcal{T}$
+    Time index of the monthly billing cycle.
+  • $n$
+    New user whose arrival triggers current optimization.
+  • $i/\mathcal{I}$
+    Set of all active users connected to the charging station, including user $n$, $\mathcal{I}=\mathcal{A}_{\text{sch}}\cup\mathcal{A}_{\text{reg}}\cup\{n\}$.
+  • $m$
+    Charging service choice selected by user $n$. $m\in\{\textit{REGULAR}{},\textit{SCHEDULED}{},\textit{LEAVE}{}\}$.
+  • $i,j/\mathcal{A}_{m}$
+    Set of existing users with service option $m$.
+  • $T^{\text{start}}$
+    Time index of current optimization, i.e. when user $n$ arrives at the station.
+  • $T_{i}^{\text{end}}$
+    The time index when user $i$ is no longer charging. This corresponds to the departure time indicated by a SCHEDULED user or the time index when a REGULAR user’s battery is full.
+  • $T^{\text{end}}$
+    Time index of latest departure across all users $i\in\mathcal{I}$. $T^{\text{end}}=\max_{i\in\mathcal{I}}T_{i}^{\text{end}}$.
+  • Parameters
+  • $\Delta t$
+    Time step of the system, in hours.
+  • $p^{\text{max}}$
+    Maximum power of a charger, in kW.
+  • $\eta$
+    Charger efficiency.
+  • $c_{\tau}$
+    Utility rate for electricity at time index $\tau$, in \$/kWh.
+  • $c_{d}$
+    Utility rate for demand charge, in \$/kW.
+  • ${\text{E}}_{i}^{\text{req}}$
+    Required energy of user $i$, in kWh.
+  • Variables
+  • $p$
+    Power profiles of all users in $\mathcal{I}$ to be optimized in current optimization, in kW.
+  • $p^{\text{sch}}_{i,\tau}$
+    Charging power for SCHEDULED user $i$ at time $\tau$ in the current optimization, in kW.
+  • $e_{i,\tau}$
+    Energy level for user $i$ at time index $\tau$, in kWh.
+  • $G_{\tau}$
+    Station charging load at time index $\tau$, in kW.
+  • $\hat{G}_{\tau\mid m}$
+    The forecasted station charging load at time index $\tau$, assuming new user $n$ chooses charging option $m$, in kW.
+  • $D_{\tau}$
+    Station peak power up to time $\tau$, in kW.
+  • $\zeta_{i}$
+    Fixed charging price for existing user $i$ at the time index of the current optimization, in \$/kWh.
+  • $z^{m}$
+    Charging tariff for new user $n$ who selects option $m$, in \$/kWh.
+  • $z$
+    Set of prices $(z^{\text{sch}},z^{\text{reg}})$ to be presented to new user $n$, in \$/kWh.
 I Introduction
@@
 We denote $\text{Pr}\left(m\mid z\right)$ to be the probability that new user $n$ chooses option $m$ when presented with the set of prices $z$.
-We model these probabilities with a two-step discrete choice model, described in Appendix A.
+We model these probabilities with a two-step discrete choice model, described in Appendix 0.A.
 Due to the non-convexity of the discrete choice model (A.5)–(A.9), optimization problem (1) is difficult to solve numerically.
@@
 Model hyperparameters are tuned with a train-validation-test split of empirical station usage data.
-More details on feature engineering and forecast implementation can be found in Appendix B.
+More details on feature engineering and forecast implementation can be found in Appendix 0.B.
 IV Numerical Study
@@
 Note that the computational complexity of the algorithms proposed in Section III is not prohibitive; all algorithms run within in a fraction of a second on a standard laptop.
-Further details on the simulation setup can be found in Appendix C.
+Further details on the simulation setup can be found in Appendix 0.C.
 IV-B Results and Discussion
@@
 Enhancing behavioral simulation, testing our algorithms for control of large aggregate loads, and experimentally deploying our control schemes with real human behavior as in [Ozturk2025, lee-adaptive] are key steps toward further improvement of joint price and power control for demand charge mitigation at workplace EV charging stations.
-Appendix A Behavioral Model Formulation
+Appendix 0.A Behavioral Model Formulation
 Our behavioral model describes the probability that user $n$ selects charging choice $m$ when presented with prices $z$.
@@
 $=\frac{e^{U_{\text{reg}}}}{e^{U_{\text{sch}}}+e^{U_{\text{reg}}}}(1-\text{Pr}(m=\textit{LEAVE}{}\mid z)).$ (A.9)
-Appendix B Time Series Forecasting Details
+Appendix 0.B Time Series Forecasting Details
 Our model forecasts the station power time series for the next $l$ timesteps.
@@
 $\forall\tau\in[T^{\text{start}},T^{\text{end}}).$
-Appendix C Simulation Setup
+Appendix 0.C Simulation Setup
 The SlrpEV dataset used to derive our simulation has a mix of REGULAR and SCHEDULED charging sessions. As we re-generate user choices using the discrete choice model, we make several assumptions: